OpenStudy (anonymous):

What is the vertical asymptote of this function: y=x/(x^2 + 1)

OpenStudy (aum):

For rational functions such as these vertical asymptotes is when the denominator goes to zero. What would make (x^2 + 1) = 0? No real values of x will make x^2 + 1 zero. So there are no vertical asymptotes.

OpenStudy (aum):

\(x^2 + 1 = 0; ~~ x^2 = -1; ~~ x = \pm i \). Only complex values of x will make the denominator zero. No real values. Therefore, no vertical asymptotes.

OpenStudy (anonymous):

Oh okay. Thanks!! ^^